# Efficient Data-Sketches and Fine-Tuning for Early Detection of Distributional Drift in Medical Imaging

### v1

@@ IV-D1 Gaussian Noise
IV-D2 Salt-and-Pepper Noise
@@ IV-D4 Poisson Noise
V Related Work
Current research [16, 17, 18, 19] sought to mitigate the limitations of data drift by utilizing transfer learning, which adapts pre-trained models to new datasets with minimal training. While this method effectively reduces the data requirements, it also exposes the models to potential inaccuracies due to noise and outliers in the data. The sensitivity of deep learning models to such imperfections can significantly impair their ability to detect subtle distributional drifts, thereby undermining the reliability of diagnostic decisions.
Most of the existing research on drift detection primarily focuses on textual datasets [20, 21, 22, 23], where distributional shifts are typically identified by monitoring changes in word embeddings or other text-based features. Drift detection in medical imaging has received relatively less attention. Unlike textual data, medical images are high-dimensional, and small, subtle changes can have significant implications. Existing techniques for drift detection in images often struggle with the complexity and high-dimensionality of image data, leading to less effective monitoring of distributional shifts. Some approaches [24, 25, 26] rely on visual inspection or traditional image processing techniques, which may not be scalable or effective for real-time detection in large image datasets.
@@ VI Conclusion
This paper introduces an advanced method for detecting distributional drift in CT-scan medical imaging by utilizing data-sketching techniques and fine-tuning a ViT model. The approach enhances anomaly detection accuracy and stability, offering a scalable solution for maintaining model performance in dynamic clinical settings. By leveraging data sketches and cosine similarity, the method enables efficient real-time image comparison, improving model robustness against subtle data shifts.

### v2

@@ V-E1 Gaussian Noise
V-E2 Salt-and-Pepper Noise
@@ VI Conclusion
This paper presents an innovative approach to detecting distributional drift in medical imaging by combining data-sketching techniques with a fine-tuned Vision Transformer (ViT) model. Our method effectively addresses challenges such as high-dimensional image data, noisy datasets, and the need for real-time detection. By leveraging data sketches for compact, noise-resilient representations and fine-tuning ViT for enhanced feature extraction, the approach achieves exceptional sensitivity to subtle data shifts, as demonstrated by its 99.11% accuracy on breast cancer CT image classification tasks. This work provides a robust and scalable solution for maintaining diagnostic model accuracy in dynamic clinical settings. Future research can explore extending this approach to other medical imaging modalities or integrating it with hospital workflows to improve practical applicability. By ensuring consistent and reliable performance, our method contributes to advancing AI-driven tools for better patient care and safety.